2.9. Let a and b be integers.

(a) Prove that the subset aZ + bZ = {ak + bl | l, k ∈ Z } is a subgroup of Z.

(b) Prove that a and b + 7a generate the subgroup aZ + bZ.


1
Expert's answer
2022-06-22T09:36:24-0400

a) Clearly "a\mathbb{Z}+b\mathbb{Z}\ne\phi". Let "ak_1+bl_1", "ak_2+bl_2" be two elements in "a\mathbb{Z}+b\mathbb{Z}," where "k_i,l_i\in\mathbb{Z}". We have "(ak_1+bl_1)-(ak_2+bl_2)=a(k_1-k_2)+b(l_1-l_2)\in a\mathbb{Z}+b\mathbb{Z}" as "k_1-k_2,l_1-l_2\in\mathbb{Z}". This implies "a\mathbb{Z}+b\mathbb{Z}" is a subgroup of "\mathbb{Z}".


b) Firstly "a,b+7a\in a\mathbb{Z}+b\mathbb{Z}". Secondly, given any "ak+bl\in a\mathbb{Z}+b\mathbb{Z}" we can write "ak+bl=a(k-7l)+(b+7a)l" , this implies "a" and "b+7a" generate "a\mathbb{Z}+b\mathbb{Z}".